T2-2 Processing costing: Cost flows (LO 2) Fill in the missing data (ignore indirect materials). Beginning balance MaterialsPlease find these numbers and explain the reasons.

Thank you.
Answer #1

a. Ending Balance of Material = Beginning inventory + Raw Material Purchased - Material issued
= 150000 + 805000 - 800000 = $155000

b. Ending WIP - Department 1 = Beginning Inventory + Material used + Conversion - Transferred to Department 2
= $60000 + 630000+516000 -1100000 = $106000

c. Material used in WIP Department 2 = Total Material issued - Material used in Department 1
= $800000 - 630000 = $170000

d. Transferred from Department 1 = $1100000

e. Transferred to Finished Goods = Beginning Department 2 + Raw Material used + Conversion + Transfer from Dep 1 - Ending Department 2
= 51000 + 170000+180000+1100000-31000 = $1470000

f. Transferred from Department 2 = $1470000

g. Ending Balance Finished Goods = Beginning Inventory + Transferred from Department 2 - Ending Inventory
= $250000 + 1470000 - 150000 = $1570000